### v1

# CERN-TH-2024-177

Abstract
We propose a novel strategy to fit experimental data using a UV complete amplitude ansatz satisfying the constraints of Analyticity, Crossing, and Unitarity. We focus on $\pi\pi$ scattering combining both experimental and lattice data. The fit strategy requires using S-matrix Bootstrap methods and non-convex Particle Swarm Optimization techniques. Using this procedure, we numerically construct a full-fledged scattering amplitude that fits the data and contains the known QCD spectrum that couples to $\pi\pi$ states below $1.4$ GeV. The amplitude constructed agrees below the two-particle threshold with the two-loop $\chi$PT prediction. Moreover, we correctly predict the $D_{2}$ phase shift, the appearance of a spin three state, and the behavior of the high-energy total cross-section. Finally, we find a genuine tetraquark resonance around 2 GeV, which we argue might be detected by looking into the decays of B mesons.
pacs:
Valid PACS appear here
I Introduction
[…]
Scattering data are measured in experiments for real energies. Yet, the physics is often hidden deep in the complex energy plane intertwined with the analytic structure of scattering amplitudes. The S-matrix program provides the tools to analytically continue measurements, and learn about the spectrum and interactions of a theory.
In this Letter, we discuss a novel approach to fit experimental scattering data using a UV complete ansatz that embodies the physical constraints of Analyticity, Crossing symmetry, and Unitarity. [^1] To the best of our knowledge, a fit ansatz satisfying all these constraints has never been constructed. We propose using non-perturbative S-matrix Bootstrap methods to construct a fit model that satisfies all those properties, with only a few tunable parameters. Those parameters are then fixed by minimizing the $\chi^{2}$ with the experimental data. We will call this strategy “Bootstrap Fit”.
We test our algorithm on $\pi\pi$ scattering data. Pions are pseudo-Goldstone bosons of the original chiral symmetry of the QCD Lagrangian. For simplicity, we assume the pions have all mass $m_{\pi}=1$, neglecting the splitting due to isospin-breaking effects. Thus, they are in the vector representation of $O(3)$ and the $2\to 2$ amplitude of $\pi^{a}\pi^{b}\to\pi^{c}\pi^{d}$ is
${\cal T}_{ab}^{cd}(s,t,u){=}{\cal A}(s|t,u)\delta_{ab}\delta^{cd}+{\cal A}(t|s,u)\delta_{a}^{c}\delta_{b}^{d}+{\cal A}(u|s,t)\delta_{a}^{d}\delta_{b}^{c}\,$ (1)
where $s,t,u$ are the Mandelstam variables and $s{+}t{+}u{=}4$. By crossing symmetry, we also have ${\cal A}(s|t,u)={\cal A}(s|u,t)$.
Figure 1: The vertical side of the boxes corresponds to the uncertainty of the mass determination, the horizontal side to $\Gamma/2$, with $\Gamma$ the width (we do not show the error on the width here). Mass and width are in the same units. Gray refers to the experimental spectrum, red to our Bootstrap estimate. The particles marked by an asterisk, are still affected by numerical systematic as explained later in section V.
Pions are the lightest states of the QCD spectrum. A two-particle $\pi\pi$ state has $G$-parity $G=+1$, and parity $P=(-1)^{J}$, where $J$ is the spin of the particle. Only resonances with the same quantum numbers can be produced in a scattering experiment. In figure 1, the experimentally observed spectrum below $1.4$ GeV is represented with gray boxes, adding only the $\rho_{3}$ above this energy. [^2] To express the spectrum in dimensionless units we use $m_{\pi}=137.3$ MeV, the average between charged and neutral pions. The vertical size of the boxes sets the error on the mass, the horizontal size represents the width of the corresponding particle (the error on the width is not shown).
The Numerical Bootstrap of pion scattering amplitudes was recently revived in Guerrieri et al. (2019) and later extended to the massless case in Guerrieri et al. (2021a). In Bose et al. (2020), the authors have refined the allowed region found in Guerrieri et al. (2019) using additional positivity constraints at low energy. By performing hypothesis testing using relative entropy, they have selected the region of low energy parameters that would match QCD low energy data and $\chi$PT. In Bose et al. (2021), they have also extracted the spectrum finding hints of emerging Regge trajectories. In He and Kruczenski (2023, 2024), the authors have used several low energy constraints below the two-particle threshold, and form factors constraints computed using perturbative QCD at energy above 2 GeV, to select a region in parameter space that nicely agrees with experimental phase shifts and low energy data. The Bootstrap of large-$N$ $\pi\pi$ amplitudes was kicked off in Albert and Rastelli (2022) and Fernandez et al. (2023), and generalized by including photons and matching with the chiral anomaly in Albert and Rastelli (2024). Finally, in Albert et al. (2024), including a minimal amount of spectrum assumptions, a Bootstrap candidate for the large-$N$ QCD amplitude was found.
Our main result is the construction of a full-fledged analytic amplitude ${\cal T}_{ab}^{cd}(s,t,u)$ that fits the experimental data. From this amplitude, we predict the spectrum and low energy parameters. In figure 1, we present in red the Bootstrap estimate of the spectrum. Both the mass and the width of the resonances match the experimental measurements, except those denoted by an asterisk that are still affected by numerical systematics. The last column has the exotic quantum numbers $2^{+}(0^{++})$. There is no such state in the PDG Navas et al. (2024). It is a genuine prediction of our Bootstrap procedure. Before discussing in detail these results, we shall first describe the two steps of our strategy.
In section II, we discuss the S-matrix Bootstrap step in constructing a candidate amplitude $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$ depending on a small set of free parameters $\Theta$ by solving a semi-definite optimization problem. Next, we construct and minimize the $\chi^{2}(\Theta)$ using a gradient-free Particle Swarm Optimization (PSO) algorithm, reviewed in section III. To evaluate the $\chi^{2}(\Theta)$ function we use the result of the Bootstrap problem $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$, and the experimental and lattice data available for the $S_{0}$, $S_{2}$, $P$, and $D_{2}$ waves, as explained in section IV, where we present the result of the Bootstrap Fit. In section V we test the amplitude obtained in different regimes by comparing it with the available experimental measurements, finding always a good qualitative agreement. Finally, in section VI we comment on possible improvements and future directions.
II Building the fit model
II.1 The analytic Bootstrap ansatz
We parametrize the pion amplitude with the $\rho$-ansatz Paulos et al. (2019); Guerrieri et al. (2019) constructed to be analytic and crossing symmetric. [^3] Here we assume maximal analyticity. See Guerrieri and Sever (2021); Elias Miro et al. (2024); Guerrieri et al. (2023a) for a complementary Bootstrap approach that uses only the rigorous analyticity domain proven by Martin Martin (1966). For our problem, it is convenient to introduce an extended version of the $\rho$-ansatz, the multi-foliation ansatz. We call foliation, a sum of the form
$\mathcal{F}^{N}_{\sigma}(s|t,u){=}\sum_{0\leq n{+}m\leq N}\alpha^{(\sigma)}_{n,m}\rho_{\sigma}(s)^{n}(\rho_{\sigma}(t)^{m}{+}\rho_{\sigma}(u)^{m})+$
$\sum_{\begin{subarray}{c}n+m\leq N\\ 1\leq n\leq m\end{subarray}}\beta^{(\sigma)}_{n,m}(\rho_{\sigma}(t)^{n}\rho_{\sigma}(u)^{m}{+}\rho_{\sigma}(u)^{n}\rho_{\sigma}(t)^{m}),$ (2)
where $\rho_{\sigma}(s)=\tfrac{\sqrt{\sigma-4}-\sqrt{4-s}}{\sqrt{\sigma-4}+\sqrt{4-s}}$ is a conformal map from the cut plane to the unit disk with center $\rho_{\sigma}(8{-}\sigma)=0$. $\mathcal{F}^{N}_{\sigma}(s|t,u)$ is manifestly symmetric in $t,u$ which automatically enforces crossing symmetry. The multi-foliation ansatz is then obtained by summing over different centers $\sigma\in\Sigma$
$\mathcal{A}^{\text{ansatz}}(s|t,u)=\sum_{\sigma\in\Sigma}\mathcal{F}^{N_{\sigma}}_{\sigma}(s|t,u).$ (3)
The intuition behind this operation is simple. A single foliation $\mathcal{F}^{N}_{\sigma}$ approximates best the amplitude in the region $|s|\approx\sigma$. For single-scale problems, tuning $\sigma$ to the desired scale is enough to achieve fast convergence Elias Miro et al. (2023); Gaikwad et al. (2024). [^4] By fast convergence we mean that there is a $\sigma$ which minimizes the difference between the true amplitude and the foliation $|\mathcal{A}-\mathcal{F}^{N}_{\sigma}|$ at fixed $N$.
[…]
II.2 The unitarity constraints
The ansatz (3) is not manifestly unitary for arbitrary values of the free parameters $\alpha^{(\sigma)}_{n,m}$ and $\beta^{(\sigma)}_{n,m}$. We impose unitarity as a numerical constraint on those coefficients. This is conveniently performed by projecting the amplitude on a set of partial waves diagonalizing the amplitude in angular momentum and flavor. First, we decompose the amplitude into the isospin channels
$\mathcal{T}^{(0)}(s,t,u)$ $=$ $3\mathcal{A}(s|t,u)+\mathcal{A}(t|s,u)+\mathcal{A}(u|s,t),$ (4)
$\mathcal{T}^{(1)}(s,t,u)$ $=$ $\mathcal{A}(t|s,u)-\mathcal{A}(u|s,t),$ (5)
$\mathcal{T}^{(2)}(s,t,u)$ $=$ $\mathcal{A}(t|s,u)+\mathcal{A}(u|s,t),$ (6)
and then project into partial waves
$t_{\ell}^{(I)}(s)=\frac{1}{32\pi}\int_{-1}^{1}d\cos\theta P_{\ell}(\cos\theta)\mathcal{T}^{(I)}(s,\theta)\,,$ (7)
where $\cos\theta=1+2t/(s-4)$. Finally, unitarity for partial waves becomes the probability conservation condition $|S_{\ell}^{(I)}|\leq 1$ for any $s>4$, any $\ell$, and $I$, with $S_{\ell}^{(I)}=1+i\sqrt{\tfrac{s-4}{s}}t_{\ell}^{(I)}$.
Experimental data show a pronounced inelasticity around $s\approx 50$ at the $K\bar{K}$ threshold where the process $\pi\pi\to K\bar{K}$ goes on-shell, especially in the $S_{0}$ channel. The probability conservation must be then replaced by the stronger condition $|S_{\ell}^{(I)}|\leq\eta_{\ell}^{(I)}$. We impose this condition as an SDP inequality of the form [^6] The effect of inelastic constraints in the S-matrix Bootstrap has been discussed in Antunes et al. (2023). The Bootstrap of full multi-particle processes has been only recently developed for branon scattering amplitudes in two space-time dimensions Guerrieri et al. (2024). Inelasticities were also obtained as an output of a Bootstrap procedure in Tourkine and Zhiboedov (2023).
$\mathcal{U}_{\ell}^{(I)}=\begin{pmatrix}\eta_{\ell}(s)+\mathop{\mathrm{Im}}S(s)_{\ell}&\mathop{\mathrm{Re}}S(s)_{\ell}\\ \mathop{\mathrm{Re}}S(s)_{\ell}&\eta_{\ell}(s)-\mathop{\mathrm{Im}}S(s)_{\ell}\end{pmatrix}\succeq 0.$ (8)
As for the functions $\eta_{\ell}^{(I)}(s)$, we use the phenomenological parametrization discussed in Garcia-Martin et al. (2011). These functions should be regarded as an additional phenomenological input. [^7] It would be interesting to consider the general mixed system of pions and kaons, and obtain the inelasticity as a result of the procedure. See Homrich et al. (2019); Bercini et al. (2020); Guerrieri et al. (2020) for S-matrix Bootstrap works on mixed-amplitudes.
II.3 Soft theorems and Spectrum Assumptions
In Guerrieri et al. (2019), it was shown that the low energy parameters of the pion amplitude, such as scattering lengths, are well inside the allowed region determined by the general S-matrix Bootstrap constraints. To restrict the allowed region, it is necessary to impose additional conditions on the amplitude. To this end, we consider two types of constraints: soft theorems, and spectrum assumptions.
Soft theorems are the consequence of spontaneously broken chiral symmetry. If pions were massless, the amplitude would vanish $\mathcal{A}\to 0$ when any of the momenta of the particles become soft. As the pions are massive, this equation is corrected by quantum effects and is no longer exact Adler (1965). Nevertheless, the existence of low energy zeroes in the partial waves is a prediction of Chiral Perturbation Theory ($\chi$PT) Weinberg (1979). For this reason, we impose that $t_{0}^{(0)}(z_{0})=0$, and $t_{0}^{(2)}(z_{2})=0$ for some $0\leq z_{0},z_{2}\leq 4$. We refer to these constraints as “chiral zeroes” conditions. [^8] At tree-level in $\chi$PT: $z_{0}=1/2$, and $z_{2}=2$.
The physical spectrum is encoded into the position of pole singularities in the second sheet of the $2\to 2$ scattering amplitude. Using the elastic unitarity condition, it is possible to relate those resonance poles to zeros of the $S$-matrix (not the amplitude!) in the first sheet. [^9] When resonances have a mass higher than multi-particle thresholds the analytic continuation becomes subtle. While resonance poles imply zeros in the first sheet, in general, zeros do not have to be the signature of a real resonance state. We impose several conditions of the form $S_{\ell}^{(I)}(s_{R})=0$, that we call “resonance zeros”, where $s_{R}$ is the complex mass squared of the particle, and $(\ell,I)$ its quantum numbers.
At this point, it is important to stress that we are agnostic about the values of the chiral and resonance zeros. In our construction, we only assume their existence with the corresponding quantum numbers. Their numerical values will be a prediction of the fit.
For the spectrum, we only assume the existence of a part of it. It turns out that to correctly reproduce the experimental data, we will need to consider at least four resonance zero constraints, one for the $\rho(770)$ in the $P$ wave, the two $f_{0}$’s in the $S_{0}$, [^10] We call the resonances $f_{0},f_{0}^{\prime},f_{0}^{\prime\prime},\dots$ to differentiate them. We will use similar notation in other channels. In the $S_{0}$ channel, we do not include the $\sigma$ in this list. and one $f_{2}$ in the $S_{2}$ wave. We will then discover the full spectrum by searching for additional dynamically generated zeros in the partial waves.
II.4 The target functional
The final step in building our fit model is choosing the Bootstrap target functional. There are in principle infinitely many options, but not all are equivalent. For our purpose, we have found it convenient to follow the objective studied in Guerrieri et al. (2019). [^11] It would be interesting to understand the relation between this objective and the regions selected in Bose et al. (2020, 2021), and He and Kruczenski (2023, 2024) using alternative principles. To explain the why, let us consider the most general ansatz (3), and fix the $S_{0}$ scattering length $a_{0}^{(0)}$ [^12] Scattering lengths are defined from the threshold expansion of the partial amplitudes $\mathop{\mathrm{Re}}t^{(I)}_{\ell}(s)=2k^{2\ell}(a_{\ell}^{(I)}+k^{2}b_{\ell}^{(I)}+\dots)$, where $k^{2}=s/4-1$ is the center of mass momentum. to a certain value, as well as for the chiral zeros $z_{0}$ and $z_{2}$. If we then look at the allowed region of the $\{a_{1}^{(1)},a_{0}^{(2)}\}$ scattering lengths we generically find a kink as in figure 2. Tuning the values of the $\{a_{0}^{(0)},z_{0},z_{2}\}$ parameters, it is possible to obtain a kink close to the experimental values of the scattering lengths. When that happens, the extremal amplitude at the kink closely resembles the pion amplitude, but not exactly. [^13] The values of scattering lengths, and subleading threshold coefficients match the experimental values, and the corresponding amplitude contains both the $\sigma$, and the $\rho$ resonance states. However, although the $\sigma$ position is compatible with the data, the width of the $\rho$ is larger than the experimental one. We have tried to fine tune the triplet $\{a_{0}^{(0)},z_{0},z_{2}\}$ in order to match the experimental values of $\sigma$ and $\rho$, but without success. We conjecture that by fine-tuning additional low energy constants it would be possible to dynamically generate the $\rho$ in the correct position. For this reason, we expect that imposing a spectrum assumption might be enough to obtain a good fit model for the pion amplitude.
Figure 2: The kink from Guerrieri et al. (2019) in the scattering length space $\{a_{0}^{(0)},a_{1}^{(1)}\}$. The allowed region is in blue. The point with error bars is the experimental determination, see also Table 2. The blue dot is the position of the kink as a result of our Bootstrap fit algorithm.
The Bootstrap problem designed to construct the fit model of the pion amplitude can be summarized as follows. {mdframed}[frametitle=Fit Ansatz,frametitlealignment=,backgroundcolor=blue!6, leftmargin=0cm, rightmargin=0cm, topline=false,bottomline=false, leftline=false, rightline=false]
Given $\Theta=\{\theta,a_{0}^{(0)},z_{0},z_{2},m_{\rho}^{2},m_{f_{0}}^{2},m_{f^{\prime}_{0}}^{2},m_{f_{2}}^{2}\}$
$\underset{\text{in }{\mathcal{A}^{\text{ansatz}}(s,t,u)}}{\text{Maximize}}$ $\text{Obj}(\theta)$
constr. by $t_{0}^{(0)}(4){=}2a_{0}^{(0)},t_{0}^{(0)}(z_{0}){=}0,t_{0}^{(2)}(z_{2}){=}0$
$S_{1}^{(1)}(m_{\rho}^{2}){=}0,S_{0}^{(0)}(m_{f_{0}}^{2}){=}0$
$S_{0}^{(0)}(m_{f_{0}^{\prime}}^{2}){=}0,S_{2}^{(0)}(m_{f_{2}}^{2}){=}0$
$s\geq 4$ $\mathcal{U}_{\ell}^{(I)}\succeq 0\,\,\text{for}\,\,\ell\in\mathbb{N},\,I=0,1,2$ (9)
The objective is the “kink functional”
$\text{Obj}(\theta)=a_{1}^{(1)}\cos\theta+a_{0}^{(2)}\sin\theta,$ (10)
which depends on the angle $\theta$ that typically is somewhere between $\pi<\theta<3\pi/2$. We let $\theta$ free, and add it to the set of fit parameters. [^14] We could consider a different objective $\text{Obj}^{\prime}(\theta,\phi)=a_{0}^{(0)}\cos\phi\sin\theta+a_{0}^{(2)}\sin\phi\sin\theta+a_{1}^{(1)}\cos\theta$, and replace the $a_{0}^{(0)}$ parameter with a new angle $\phi$. This would lead to an equivalent formulation of the Bootstrap ansatz (9). From this view, it is evident we are moving along the two-dimensional boundary of the three-dimensional parameter space of the scattering lengths. This can be generalized by adding an arbitrary number of low energy observables, which amounts to scanning over a larger space of amplitudes.
The quantities collectively denoted by $\Theta$ are the parameters of the fit and the input for the optimization problem (9). We solve this problem using the standard SDPB solver Simmons-Duffin (2015); Landry and Simmons-Duffin (2019). The corresponding extremal amplitude depends on the choice of $\Theta$. Next, we will explain how to optimize on $\Theta$. For this problem, the size of the parameter set is $|\Theta|=12$.
III Gradient free optimization and particle swarm
For any choice of $\Theta$, the optimal solution of the Bootstrap problem (9) yields a model for the scattering amplitude $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$. To choose $\Theta$, we construct the $\chi^{2}(\Theta)$ using the Bootstrap amplitude and the experimental data, and we minimize it. The dependence of the model $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$ is non-linear in $\Theta$, hence we expect the $\chi^{2}(\Theta)$ to be a non-convex function.
In this Letter, we explore an algorithm especially suited to this class of problems, the Particle Swarm Optimization algorithm (PSO) Shi and Eberhart (1998). [^15] We are grateful to Balt van Rees for pointing out this method. The PSO is a standard algorithm designed for solving non-convex problems, with a wide range of applications – see Clerc (2006) for an introduction. The PSO is also gradient-free and does not require the computation of derivatives of the Bootstrap solution in $\Theta$ [^16] As explained in Reehorst et al. (2021), it is possible to efficiently compute the gradient of a Bootstrap problem solved with the interior point method as the one implemented in SDPB.
The general idea is simple, but the implementation details may depend on the problem. We start with $n_{p}$ particles at random positions $\Theta^{(i)}_{0}$ to which we assign random velocities $v^{(i)}_{0}$. Positions are drawn from a distribution peaked in the region where we believe the fit parameters live. Velocities are of the same order as the size of this region.
At step $n$, we update their positions following the rule
$v^{(i)}_{n+1}$ $=$ $\omega v^{(i)}_{n}+c_{1}r_{1}(\Theta^{(i)}_{n}-X^{(i)}_{n})+c_{2}r_{2}(\Theta^{(i)}_{n}-Y_{n}),$
$\Theta^{(i)}_{n+1}$ $=$ $\Theta^{(i)}_{n}+v^{(i)}_{n+1}.$ (11)
We denote with $X_{n}^{(i)}$ the position with the best $\chi^{2}$ found by the $i$-th particle along its trajectory up to the step $n$ and with $Y_{n}$ the position with the best $\chi^{2}$ found by the entire swarm. $r_{1,2}$ are random numbers uniformly distributed between $[0,1]$. To construct each step, we first evaluate the $\chi^{2}(\Theta_{n}^{(i)})$ for each particle, then we use this information to update the $X_{n}^{(i)}$, and $Y_{n}$.
Three parameters control the algorithm performance: the inertia $\omega$, the cognitive coefficient $c_{1}$, and the cooperation coefficient $c_{2}$. The convergence property of the algorithm depends on the choice of these three coefficients. [^17] In other versions of the algorithm, those parameters can be promoted to be dynamical. The details of our implementation of the PSO algorithm, which contains an additional adaptive velocity prescription as in Xu (2013), can be found in Appendix B.
The algorithm runs for $N_{\text{iter}}$ steps. The best position found by the swarm at the last step $Y_{N_{\text{iter}}}$ is our estimate for the best fit parameters which gives the minimum value of $\chi^{2}(\Theta)$.
IV Bootstrap Fit results
Figure 3: The four channels used to fit the pion amplitude. The points respectively in orange and green are experimental and lattice data. The thick blue curve is the best fit, the light blue cloud is given by all the curves with sub-optimal $\chi^{2}$. (Notation: $S_{0}$ stands for the $(I,\ell)=(0,0)$ channel, $P$ for the $(1,1)$, $S_{2}$ for the $(2,0)$, and $D_{0}$ for the $(0,2)$.
We construct the $\chi^{2}(\Theta)$
$\chi^{2}(\Theta)=\sum_{i\in\text{data set}}\left(\frac{\delta_{i}^{\text{exp}}-\delta_{\Theta}^{\text{Ansatz}}(s_{i})}{\Delta\delta_{i}^{\text{exp}}}\right)^{2}$ (12)
using experimental and lattice phase shifts $\delta_{i}^{\text{exp}}$, where the index $i$ stands collectively for the quantum numbers $(I,\ell)=\{(0,0),(1,1),(2,0),(0,2)\}$, and the energy $s_{i}$ of the measurement. Our ansatz for the phase shift is obtained by projecting $\mathcal{A}^{\text{ansatz}}_{\Theta}$ in partial waves and using the definition $S_{\Theta}=\eta_{\Theta}\,\text{exp}(2i\delta_{\Theta})$.
The input data used in (12) are from Protopopescu et al. (1973); Estabrooks and Martin (1974); Grayer et al. (1974); Losty et al. (1974); Hoogland et al. (1977), except for the $S_{0}$ wave. The data in the $S_{0}$ channel above $s\approx 20$ extracted from the old experiments are often incompatible and suffer from unknown systematic errors. Following Pelaez and Yndurain (2005), we used only a selection of data points. [^18] We are grateful to Jose Ramon Pelaez for highlighting this point. Close to the threshold, the experimental situation was cleared by the CERN experiment NA48/2 Batley et al. (2010). We also include lattice data from the RBC and UKQCD collaborations in the $S_{0}$ and $S_{2}$ channels extrapolated at the physical pion mass Blum et al. (2021). [^19] For the lattice data points, the errors on the phase shift and the energy are strongly correlated. We do not include this effect in this work. In this case, we use the isospin symmetric value of the pion mass $m_{\pi}=135$ MeV to express energies in dimensionless units.
The result of the Bootstrap fit is shown in figure 3. The orange data points are experimental, the green are taken from the lattice. The dark blue curve is the best fit with minimum $\chi^{2}(\Theta)\approx 40$. We do not assign a statistical meaning to the value of the $\chi^{2}$, but take it as a likelihood measure. The largest contribution to the $\chi^{2}$ comes from the $P$ wave, with $\chi^{2}\approx 30$, where the experimental error is almost negligible. We plan to investigate this channel with more care in the future, replacing the phase shift data with the determination from the pion form factors de Troconiz and Yndurain (2005), and with lattice extrapolations Boyle et al. (2024). The second largest contribution of order ten is due to the $S_{0}$ wave. Here we observe a systematic error in reproducing the phase above the $K\bar{K}$ threshold $s\approx 50$ (the $\chi^{2}$ below this energy is of order one). We believe this is due to the lack of an explicit threshold in our ansatz at that scale, as corroborated by the slower convergence of the Bootstrap problem, see also Appendix A. Both $S_{2}$ and $D_{0}$ waves have $\chi^{2}$ of order one.
We run the PSO algorithm for $N_{\text{iter}}=60$ steps, using $n_{p}=10$ particles. We explore 600 points in the $\Theta$ parameter space and at each point, we construct the full analytic amplitude solving the Bootstrap problem (9). To estimate errors we consider a subset of amplitudes, and compute the weighted average using the value of the $\chi^{2}$. We observe that 50% of our set have $\chi^{2}<100$, and 15% have $\chi^{2}<50$. We compare the errors estimated using these two cutoffs and find that those obtained with the looser one are a factor of two larger. The light-blue curves in figure 3 are drawn from amplitudes with $\chi^{2}(\Theta)<50$. In the remainder of this paper, we will follow the same color scheme.
In the inset in figure 3, we zoom in on the NA48/2 and lattice data close to the threshold. The point at the kaon mass $s\approx 12$ is from lattice Blum et al. (2021). This point is crucial to help stabilize the amplitude between the threshold and the cluster of points around $s\approx 40$.
$\Theta$ Bootstrap Fit Literature
$a_{0}^{(0)}$ $0.217\pm 0.002$ $0.2196\pm 0.0034$ Batley et al. (2010)
$z_{0}$ $0.368\pm 0.008$
$z_{2}$ $2.040\pm 0.004$
$m_{\rho}$ (5.541–5.546)+$i$(1.04-1.05) (5.54–5.57)+$i$(1.03–1.08)
$m_{f_{0}}$ (7.17–7.24)+$i$(0.29–0.39) (7.14–7.36)+$i$(0.29–0.51)
$m_{f_{0}^{\prime}}$ (9.54–9.76)+$i$(2–3.2) (9.1–10.5)+$i$(0.9–4.3)
$m_{f_{2}}$ (9.28–9.30)+$i$(1.43–1.51) (9.18–9.34)+$i$(1.31–1.61)
$\theta/\pi$ $1.328\pm 0.026$
Table 1: Our estimate of the fit parameter compared and corresponding values quoted in the literature. Experimental values of the resonances are given by the PDG average Navas et al. (2024).
In Table 1, we list our estimate for the fit parameters and the corresponding errors. The error in this table is conservative and estimated using all the points with $\chi^{2}(\Theta)<100$ corresponding to half of the whole dataset of amplitudes produced.
Beyond the statistical error, there are three more sources of systematics. The first concerns the convergence of the Bootstrap model $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$ in equation (3) that depends on $N_{\text{vars}}$ the number of free variables $\alpha_{n,m}^{(\sigma)}$, and $\beta_{n,m}^{(\sigma)}$. We used two ansatzes respectively with $N_{\text{vars}}=397$ and $N_{\text{vars}}=547$. All the results we quote in this paper are taken from the latter. The difference between the fit parameters estimated from the two ansatzes is comparable with the statistical error unless stated otherwise. The second source comes from the PSO search algorithm, which does not guarantee finding the global minimum. We checked the stability of our fit by performing different searches varying the search parameters that produced almost the same result. The last source of systematics comes from the choice of the Bootstrap functional. In this case, we have tested an alternative Bootstrap formulation obtaining the same fit.
V Predictions
The big advantage of our strategy is that the result of the fit is a full analytic amplitude. Next, we study its properties away from the region directly constrained by the fit.
V.1 Amplitude below threshold
Figure 4: Partial wave amplitudes $t_{\ell}^{(I)}$ as function of $s/m_{\pi}^{2}$ in the real strip $40<s<$ 4.
The $\chi$PT expansion is a reliable approximation of the pion amplitude at low energy. In figure 4, we plot the partial waves $t_{0}^{(0)}$, $t_{0}^{(2)}$, and $t_{1}^{(1)}$ of our best fit (the solid black line) in the sub-threshold region $40<s<$ 4, and we compare it with perturbation theory Bijnens et al. (1996). The tree level, denoted with the dashed green line is given by
$t_{0}^{(0)}=\frac{2s-1}{32\pi f_{\pi}^{2}},\quad t_{1}^{(1)}=\frac{s-4}{96\pi f_{\pi}^{2}},\quad t_{0}^{(2)}=\frac{2-s}{16\pi f_{\pi}^{2}}\,,$ (13)
where $f_{\pi}$ is the pion decay constant. The higher loop expressions are more involved and are also plotted in figure 4. The two chiral zeros $z_{0}$, and $z_{2}$ are also visible in the figure. We emphasize that the only input in this region is the existence of the two chiral zeros, not their position.
Next, we extract the threshold parameters from our amplitude and compare them with the values of Colangelo et al. (2001) in Table 2. We find a very nice agreement for the leading and sub-leading threshold coefficients for all waves below $\ell=2$. The scattering lengths we extract are different for higher spins but of the same order. It would be interesting to investigate the reason for this discrepancy.
Bootstrap Fit Literature
$a_{0}^{(2)}$ $(-0.432\pm 0.001)\times 10^{-1}$ $(-0.444\pm 0.012)\times 10^{-1}$
$a_{1}^{(1)}$ $(0.380\pm 0.002)\times 10^{-1}$ $(0.379\pm 0.05)\times 10^{-1}$
$b_{0}^{(0)}$ $0.265\pm 0.030$ $0.276\pm 0.006$
$b_{0}^{(2)}$ $(-0.797\pm 0.002)\times 10^{-1}$ $(-0.803\pm 0.012)\times 10^{-1}$
$b_{1}^{(1)}$ $(0.61\pm 0.02)\times 10^{-2}$ $(0.57\pm 0.01)\times 10^{-2}$
$a_{2}^{(0)}$ $(0.53\pm 0.11)\times 10^{-2}$ $(0.175\pm 0.003)\times 10^{-2}$
$a_{2}^{(2)}$ $(0.51\pm 0.18)\times 10^{-3}$ $(0.170\pm 0.013)\times 10^{-3}$
$a_{1}^{(3)}$ $(1.5\pm 0.4)\times 10^{-4}$ $(0.56\pm 0.02)\times 10^{-4}$
Table 2: Except for the values of $a_{0}^{(2)}$ taken from Batley et al. (2010), all other threshold parameters are taken from Colangelo et al. (2001).
V.2 The $D_{2}$ partial wave
Figure 5: Top figure, elasticity for the $D_{2}$ wave as function of $s/m_{\pi}^{2}$. The black line is the parametrization used in Pelaez and Yndurain (2005). Bottom figure, phase shift for the $D_{2}$ wave, and comparison with the available experimental data. Green and orange points are the two possible determinations for the phase taken from Hoogland et al. (1977).
In figure 5, we plot the elasticity $|S_{2}^{(2)}|$ and the $\delta_{2}^{(2)}$ phase shift prediction. The color scheme follows the one defined in figure 3. The data for the phase shift are taken from Hoogland et al. (1977), orange and green correspond to two possible determinations. We compare the elasticity profile with the phenomenological parametrization from Pelaez and Yndurain (2005). It is interesting to notice a dip in the unitarity around $s\approx 90$. The dip is insensitive to Bootstrap systematics. We think that the emergent particle production in this channel is a consequence of the inelasticity profile injected in the $D_{0}$ wave. This effect is compatible with our expectation that including mixed amplitudes with kaons and pions might lead to a correct prediction for the inelasticity.
V.3 Spectrum
Beyond the spectrum assumed and fitted using the experimental data, we also observe several zeros dynamically generated in our construction.
We begin by discussing the $f_{0}(500)$ resonance, commonly referred to as the $\sigma$. We summarize the phenomenological determinations of its position in figure 6. The pink ellipses denote all the estimates of its complex mass reported in the PDG since 2001 Navas et al. (2024). The blue ellipse is the PDG average. Our estimate is represented by the black ellipse highlighted in the inset.
Figure 6: Determinations of real $\mathop{\mathrm{Re}}m_{\sigma}$ and imaginary part $\mathop{\mathrm{Im}}m_{\sigma}$ of the mass of the $\sigma$. The black ellipse highlighted in the inset is our Bootstrap determination.
To estimate the position of the $\sigma$ we look for a zero at low energy in the $S_{0}^{(0)}$ $S$-matrix using the Newton method. We repeat the operations for all amplitudes with $\chi^{2}(\Theta)<50$, and perform a weighted average to predict the mass and determine the error. The red dots in the inset are the individual determinations from this sample. Our final estimate is reported in Table 3. In figure 7, we also show the density plot for $|S_{0}^{(0)}|$ in the upper-half complex $s$ plane. The left-most zero is the $\sigma$. In the same plot, we observe two additional resonances that we identify as the $f_{0}(980)$, and $f_{0}(1370)$.
Our $\sigma$ determination is shifted away from the center of the PDG average. This is correlated with the lattice point at $s=m_{K}^{2}\approx 12$ which was not used in previous studies, and impacts the growth of the phase in the $S_{0}$ wave.
Prediction PDG average
$m_{\sigma}$ (2.97–3.01)+$i$(2.07–2.11) (2.9–5.8)+$i$(0.7–5.8)
$m_{\rho^{\prime}}$ (8.9–10.1)+$i$(0.5–0.6) (10.4–11)+$i$(1.4–1.5)
$m_{\rho_{3}}$ (16.1–17.3)+$i$(0.51–0.85) (12.2–12.4)+$i$(0.63–0.73)
$m_{\mathbb{T}}$ (14.3–14.9)+$i$(2.3–2.5) ?
Table 3: Experimental values of the resonances are taken from Navas et al. (2024).
Figure 7: $|S_{0}^{(0)}|$ in the upper half complex $s$ plane plotted for the best-fit amplitude. We highlight with black dots the three scalar resonances found in this channel.
We also find an additional state in the $P$-wave. The density plot for the $|S_{1}^{(1)}|$ in the upper half $s$ plane is shown in figure 8 for the best-fit amplitude. The two black dots are zeros of $S_{1}^{((1)}$. The left-most zero is identified as the $\rho(770)$ resonance. Its position is fixed by the fitting procedure. The second black dot on the right is dynamically generated and has a mass 5% lower than the experimental determination of the $\rho(1450)$. However, its width is half of the experimentally measured value. This resonance still suffers from systematics: we do not have the inelasticity profile at this energy, and $N_{\text{vars}}$ of the Bootstrap problem is not large enough to accurately describe this region. We suspect that dealing with this systematics might lead to the correct width in our procedure.
Figure 8: $|S_{1}^{(1)}|$ density profile in the upper half complex $s$ plane. We highlight with black dots the two $\rho$ resonances in our amplitude.
For higher spins, our amplitude dynamically generates a spin $\ell=3$ resonance, this is the first sign of Reggeization. This is typical in the amplitudes constructed using the S-matrix Bootstrap where the high energy behavior instead of being erratic is physical, although it converges slowly, see also section V.5. The mass of the $\rho_{3}$ obtained is $20$% above the experimental value, and its width is larger. This resonance, however, is still affected by the convergence of the Bootstrap ansatz, and improving the numerics might lead to a better agreement. [^20] The other possibility is that to obtain the precise Regge trajectory we might need additional fine-tuning, which can be realized by enlarging the set of fit parameters or generalizing the Bootstrap problem we use to construct the amplitude.
V.4 The tetraquark
Looking at the $S_{2}$ wave, we find a surprise. If we continue to high energy, figure 9, we observe a broad resonance with isospin $I=2$. Such a state would be a genuine tetraquark. In the real world, if we reintroduce the electric charge, such particles should also carry a charge two. Its mass and width in physical units are around 2 GeV and 600 MeV. It might be interesting to look at the invariant mass distributions of the $\pi^{\pm}\pi^{\pm}$ pairs in decays that leave enough energy to reach the 2 GeV scale, e.g. $B^{+}\to(D^{-}/K^{-},/\pi^{-})\pi^{+}\pi^{+}$ or the charge conjugate ones.
Curiously, we know an example in two dimensions where this scenario is qualitatively realized. This is the two-dimensional theory of the QCD flux-tube. There, the role of the QCD $\rho$ is played by the axion, and the role of the $\sigma$ is played by the dilaton. In this model, the analogy is striking because we have a sharp axion and a broad dilaton Gaikwad et al. (2024). In Elias Miró et al. (2019a) it was noticed that crossing symmetry enforces a broad resonance in the symmetric channel. A similar mechanism might be responsible for the presence of the tetraquark observed in our amplitude.
This zero does not suffer from Bootstrap systematics. However, we have not imposed any inelasticity in this region, and inelastic effects might slightly modify the position of this resonance. It would be interesting to test this hypothesis and make our prediction stronger.
Figure 9: High energy $S_{2}$ partial wave phase shift profile. In blue the best fit, in light blue all amplitudes with suboptimal $\chi^{2}$. We see a jump in the phase around $s\approx 200$ that signals the presence of a resonance. In the inset, the parameters are extracted from the position of the zero.
V.5 High Energy
Figure 10: Total cross sections for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}\to\pi^{-}\pi^{-}$. In blue the best fit for $N_{\text{vars}}=547$, in light blue the best fit for $N_{\text{vars}}=397$. Above 2 GeV there is still room to improve numerics, but there is already a very good agreement between Bootstrap and the experimental data Biswas et al. (1967); Robertson et al. (1973); Cohen et al. (1973); Losty et al. (1974); Hanlon et al. (1976); Abramowicz et al. (1980); Zakharov and Sergeev (1984).
Finally, we discuss the Regge behavior of our amplitude and compare it with experimental data.
The emergent high-energy behavior of S-matrix Bootstrap amplitudes is the most surprising aspect recently understood. Although the Bootstrap ansatz in (3) goes to a constant at infinity, there is an intermediate regime in which the numerical amplitudes behave according to Regge theory, showing at large $s$ and fixed $t$ the expected behavior $\mathcal{A}\sim s^{\alpha(t)}$ Bhat et al. (2024); Gumus (2023); Elias Miró et al. ; Correia et al. .
In figure 10, we compare the experimental total cross sections for $\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}$ with those extracted from our best-fit amplitude in blue [^21] In isospin components $\sigma_{\pi^{+}\pi^{-}}\propto\tfrac{1}{6}(2\mathcal{T}^{(0)}+3\mathcal{T}^{(1)}+\mathcal{T}^{(2)})$, and $\sigma_{\pi^{-}\pi^{-}}\propto\mathcal{T}^{(2)}$.
$\sigma_{\pi^{+}\pi^{-}}(s)$ $=$ $\frac{\mathop{\mathrm{Im}}\left[\mathcal{A}(s|t,u)+\mathcal{A}(t|s,u)\right]}{\sqrt{s(s{-}4)}}\biggr{|}_{t=0},$
$\sigma^{\pi^{-}\pi^{-}}(s)$ $=$ $\frac{\mathop{\mathrm{Im}}\left[\mathcal{A}(t|s,u)+\mathcal{A}(u|s,t)\right]}{\sqrt{s(s-4)}}\biggr{|}_{t=0},$ (14)
The solid blue curve is from our fit with the highest $N_{\text{vars}}=547$, while the light blue is the best fit with $N_{\text{vars}}=397$.
In the $\sigma_{\pi^{+}\pi^{-}}$ we can observe nice peak structures corresponding to the various resonances in the spectrum (the first and highest is the $\rho$ peak). In the $\sigma_{\pi^{-}\pi^{-}}$ we do not expect any peak due to resonances, except for the tetraquark around 2 GeV. At asymptotically high energies, we expect a constant behavior compatible with the pomeron exchange. As we increase $N_{\text{vars}}$ we expect the Bootstrap amplitude to better approximate this behavior which is not automatically imposed in our ansatz (3).
VI Discussion and outlook
This Letter proposes a novel constrained approach to fit experimental data. From our perspective, the advantage we foresee compared to other methods is the control of the theoretical approximations. The amplitude model used to fit the data is a genuine analytic and crossing symmetric function compatible with non-perturbative unitarity. The algorithm mixes semi-definite and non-convex optimization techniques: for the Bootstrap part we use SDPB Simmons-Duffin (2015); Landry and Simmons-Duffin (2019), and for the PSO we have coded a simple Mathematica notebook.
The $\pi\pi\to\pi\pi$ amplitude constructed has many interesting features. It reproduces the experimental data, correctly predicts the whole $G$-parity plus spectrum below 1.4 GeV, agrees with the $\chi$PT at low energy, and with the high-energy regime. We correctly predict an additional resonance heavier than the $\rho$ in the $P$ wave with a comparable mass to the real world $\rho(1450)$, and a spin three resonance 20% heavier than the experimental result. The $D_{2}$ phase shift and inelasticity agree with experimental data and phenomenological models.
Figure 11: Summary of the bootstrap fit procedure.
Surprisingly, we predict a tetraquark state with a mass of around 2 GeV in the $I=2$ channel. This is a nonperturbative prediction, and we are not aware of any other model making the same statement. Given that the amplitude constructed correctly describes and predicts several key features of $\pi\pi$ scattering amplitudes, we are quite confident this particle should be there, and it would be interesting to look for it in the decays of $B$ mesons.
There are many improvements we can implement to obtain a better model for the pion amplitude.
First, it would be interesting to include the correlation matrix for the lattice data Blum et al. (2021) in the $S_{0}$ wave, and investigate their effect on our determination of the $f_{0}(500)$ mass parameters. For the $P$ we could also include the $\delta_{1}^{(1)}$ phase extracted from form factors de Troconiz and Yndurain (2005) and the lattice QCD extrapolation at the physical pion mass Boyle et al. (2024).
The kink functional (10) used to construct our fit model may seem ad hoc and not easy to generalize to other physical systems. For this reason, we plan to repeat the fit using different optimization problems Guerrieri et al. . The existence of alternative constructions can also be considered a source of systematic that should be included in our estimates, which we currently do not account for. [^22] We have already tested a different fit approach which gives almost identical results.
We believe replacing the spectrum assumption and generating it dynamically would be ideal. This could be done by optimizing larger linear combinations of low-energy constants, generalizing the kink functional in equation (10), as it was already done in several S-matrix Bootstrap works Elias Miró et al. (2019b); Guerrieri et al. (2021a); Bose et al. (2021); Guerrieri et al. (2021b, 2023b); Häring et al. (2024); Acanfora et al. (2024).
It would also be interesting to analytically continue the best-fit solution in spin, as done in Acanfora et al. (2024), and examine the complexified Regge trajectories. This would require running with a larger ansatz to have enough numerical precision. We could investigate whether the three $0^{+}(0^{++})$ scalar resonances below 1.4 GeV are subtraction poles or part of Regge trajectories.
Another systematic of our fit procedure is the input of the shape of the inelasticity. To make the fit self-consistent we could enlarge the set of amplitudes including mixed processes involving pions and kaons. By doing so, we would have access to many additional resonances, and we could use more experimental data as input. The inelasticity profile would be an output of the fit.
It would also be interesting to use our amplitude to measure form factors, see Karateev et al. (2020); Correia et al. (2023); Cordova et al. (2024). Fixing our parametrization, we could read off the form factors as an output, and check consistency with high-energy sum rules recently discussed in He and Kruczenski (2023, 2024) for the case of $\pi\pi$ scattering. [^23] Knowledge of the form factors computed using other methods can also be used to constrain the amplitude, see e.g. Chen et al. (2022)., where the authors used Hamiltonian truncation to compute them.
Ultimately, how can we extend our method and determine rigorous bounds on the position of the resonances in figure 1? It would be interesting in this sense to generalize the strategy proposed in Gabai and Yin (2022), where the authors obtained bounds on the position of resonances in the complex $s$ plane in the case of Ising Field Theory in $1+1$ dimensions.
Finally, we could explore processes involving other particles rather than pions. For example, Glueball couplings determine the dependence of the lattice energy levels on the Glueball-Glueball phase shift in a non-perturbative way. Combining the scattering amplitudes constructed in Guerrieri et al. (2023a) with the lattice energy levels we could learn about observables that may seem hopeless to measure.
Acknowledgements
We thank Mattia Bruno, Gilberto Colangelo, Miguel Correia, Barak Gabai, Victor Gorbenko, David Gross, Aditya Hebbar, Denis Karateev, Harish Murali, Jose Ramon Pelaez, Joao Penedones, Alessandro Pilloni, Jiaxin Qiao, Riccardo Rattazzi, Slava Rychkov, Balt van Rees, David Simmons-Duffin, Pedro Vieira, and Alexander Zhiboedov for useful discussions.
We thank Joao Penedones, Alessandro Pilloni, Balt van Rees, Pedro Vieira, Mark Wise and Alexander Zhiboedov for the many useful comments on the draft.
Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI. AG is supported by the European Union - NextGenerationEU, under the programme Seal of Excellence@UNIPD, project acronym CluEs. The work of KH is supported by the Simons Foundation grant 488649 (Simons Collaboration on the Nonperturbative Bootstrap), by the Swiss National Science Foundation through the project 200020 197160, through the National Centre of Competence in Research SwissMAP and by the Simons Collaboration on Celestial Holography. NS’s work is supported by Simons Foundation grant 488657 (Simons Collaboration on the Nonperturbative Bootstrap).
Figure 12: Real (blue), imaginary (red), and absolute value (green) of the phase shifts for various channels. Thick lines are from our best numerics with $(N,M)=(14,12)$, light lines for $(N,M)=(12,10)$. The inelasticity profiles used as a constraint are denoted in black. We do not draw it for the channels where we impose the simple probability conservation $|S_{J}^{(I)}|\leq 1$.
Appendix A Details of the S-matrix bootstrap numerics
The ansatz used to produce the results of this letter has multi-foliation $\Sigma=\{20/3,30,50,86\}$. As discussed in Gaikwad et al. (2024), to best approximate resonant structures it is convenient to set the centers of the foliations around their expected positions. That explains our choice for the set $\Sigma$ which is tuned on the expected position of the $\rho(770)$ with $m^{2}_{\rho}\approx 30$, the $f_{0}(980)$ with $m^{2}_{f_{0}}\approx 50$, and the $f_{2}(1260)$ with $m^{2}_{f_{2}}\approx 86$.
Let us call $N=N_{20/3}$, and $M=N_{\sigma}$, when $\sigma$ is any other foliation. To produce the plots in the main text we used $N=14$, and $M=12$ for a total of $N_{\text{vars}}=547$ variables. We have also performed a second run with $N=12$, and $M=10$ for a total of $N_{\text{vars}}=397$ to check the systematic due to the size of the ansatz.
To impose the unitarity constraints we project numerically our ansatz into partial waves. We use a union of grids, one for each foliation. The grid is defined through the map
$s(\phi)=\frac{\sigma-(8-\sigma)\cos\phi}{1+\cos\phi}$ (15)
where $0\leq\phi\leq\pi$ is the upper boundary of the unit disk. We finally discretize $\phi$ on the roots of the Chebyschev polynomials
$\phi_{k}=\frac{\pi}{2}\left(1+\cos\frac{\pi k}{N_{\text{points}}+1}\right),$ (16)
where $k=1,\dots,N_{\text{points}}$. The maximum number of points used is $N_{\text{points}}=300$ for $\sigma=20/3$, and $N_{\text{points}}=150$ for all other foliations. We have performed numerical tests using grids with different numbers of points finding no significant dependence on it.
We have projected our ansatz onto partial waves (7) up to spin $\ell=12$, and run our numerics with both maximum spin $L_{\text{max}}=10,12$. The reason that we can keep $L_{\text{max}}$ so low is due to the addition of improved positivity constraints Elias Miro et al. (2023)
$\text{Im}\,\mathcal{T}^{(I)}(s,t)-16\pi\sum_{\ell=0}^{L_{\text{max}}}(2\ell+1)P_{\ell}(1+2\tfrac{t}{s-4})\text{Im}\,t_{\ell}^{(I)}(s)\geq 0$ (17)
for $40\leq t<$ 4, $s>4$, and $I=0,1,2$. By imposing this condition we constrain the tail of spins higher than $L_{\text{max}}$ that we do not explicitly bound with numerical unitarity.
In figure 12 we show the partial S-matrices $S_{J}^{(I)}$ obtained using two ansatzes respectively with $(N,M)=(14,2)$ (dark lines), $(N,M)=(12,10)$ (light lines). We plot their real (blue), imaginary (red), and absolute values (green). The black solid line is the inelasticity profile due to $\pi\pi\to K\bar{K}$ taken from Garcia-Martin et al. (2011) (for the $S_{0}$ channel we assume the big-dip scenario) that we impose up to $s\approx 100$. Above that energy we keep the inequality $|S_{J}^{(I)}|\leq 1$.
The unitarity constraints are sufficiently saturated in all these channels except for the region around $s\approx 100$ in the P wave around the position of the $\rho^{\prime}$, and in the $F$ wave around $s\approx 250$ at the position of the $\rho_{3}$. These dips in unitarity around resonances are a typical numerical artifact of our finite $N_{\text{vars}}$ truncation. It would be interesting to increase both $N$ and $M$ to higher values and see if the positions of these particles better align with their experimental determinations.
Figure 13: History of the PSO search in the plane $(z_{0},z_{2})$. At each step $n$, we plot in black the positions of the particles and highlight the best individualistic positions $X_{n}^{(i)}$ in blue, and the collective best position $Y_{n}$ in red. As $n$ increases the particles slowly concentrate around the optimal region. The complete movie can be seen in https://giphy.com/gifs/Q5mtDePqPtfSqaQI2z.
Appendix B Particle Swarm implementation
Here we describe our implementation of the PSO algorithm in detail.
To perform the search we first choose a region in the fit parameter space where we expect to find the minimum of the $\chi^{2}(\Theta)$ function. In our case, we choose a region located around the phenomenological estimates of the fit parameters. The position of each particle is initialized with random numbers drawn from a Gaussian distribution, with the mean at the center of the region and variance proportional to its size, which we denote by $\vec{\epsilon}$. As for the number of particles, we use $n_{\text{p}}=10$. We do a similar operation for the initial velocities of the particles: we pick them from a Gaussian distribution with zero mean and variance $\vec{\epsilon}$.
Then we start the algorithm by updating the position and velocities of the particles according to (11). We also enforce the rule that when a particle is sent outside of the box by the iteration step, we place it back on the wall with the opposite velocity.
To efficiently minimize the $\chi^{2}(\Theta)$ we use a small variation of the standard PSO called “adaptive PSO” which uses the velocity information to tune the value of $\omega$, as described in Xu (2013). We use $c_{1}=c_{2}=1.4962$, and vary $\omega$ ranging from $0.3\leq\omega\leq 0.8$ starting from the highest value $0.8$, and then decrease it or increase it at each step depending on the value of the average velocity of the swarm.
Without the adaptive step, the swarm will run until it reaches an agreement on the best position found and eventually will start exploring the neighbor of the optimal one. Since, in our case, evaluating $\chi^{2}(\Theta)$ is as computationally expensive as an entire S-matrix Bootstrap run, we adopt a strategy to reduce the number of steps required for convergence.
The idea is to tune the friction parameter $\omega$ according to some control function. The friction $\omega$ controls how well the swarm will explore the parameter space. If $\omega$ were of order one, the particles would wildly explore the space with the risk of never agreeing on the minimum. If $\omega$ were too small, the particles would immediately freeze without exploring much.
In Xu (2013), it is suggested to look at the total velocity of the swarm at each step $n$ defined as
$\mathbb{V}_{n}=\frac{1}{n_{\text{p}}|\Theta|}\sum_{i=1}^{n_{\text{p}}}|v_{n}^{(i)}|,$ (18)
and compare it with a proposed ideal velocity profile
$\mathbb{V}^{\text{ideal}}_{n}=\frac{|\epsilon|}{2}\left(1+\cos\left(\frac{\pi n}{0.95N_{\text{iter}}}\right)\right).$ (19)
If $\mathbb{V}_{n-1}\geq\mathbb{V}^{\text{ideal}}_{n}$, we set $\omega_{n}=\max(\omega_{n-1}-\Delta\omega,\omega_{\text{min}})$, otherwise $\omega_{n}=\min(\omega_{n-1}+\Delta\omega,\omega_{\text{max}})$. In our case, $\omega_{\text{max}}=0.8$, $\omega_{\text{min}}=0.3$, and $\Delta\omega=0.1$.
The expression of $\mathbb{V}^{\text{ideal}}$ contains the parameter $N_{\text{iter}}$ which is the max number of iterations of the search. The ideal velocity is a monotonic decreasing function that starts at $|\epsilon|$ when $n=0$. As we approach the end of the run when $n\sim N_{\text{iter}}$, the ideal velocity is so small that it forces the friction to decrease to its minimum value $\omega_{\text{min}}$ hence freezing the search in a neighbor of the global best found up to that point.
The danger of this approach is that if we choose $N_{\text{iter}}$ too small, we might end up in a local minimum. For this reason, we performed two searches with different $N_{\text{iter}}=30,60$.
In figure 13 we plot the positions of the particles in the $(z_{0},z_{2})$ plane at different steps $n$ in the run with $N_{\text{iter}}=60$. The black dots correspond to the actual positions of the particles at step $n$. The blue dots correspond to each particle’s best position up to step $n$, and the red dot is the best found by the swarm at the same step. At $n=0$, black and blue dots coincide. The gray area is the box where we perform the search.